OpenStudy (anonymous):

Three horses A,B and C are in race. A is twice likely to win as B and B is twice as to win as C.What is the probability that B or C wins?

OpenStudy (kropot72):

Let P(A) = x Then we can write: P(B) = x/2 and P(C) = x/4 Assuming that dead heats are not allowed, the sum of the probabilities must equal 1: \[\large x+\frac{x}{2}+\frac{x}{4}=1\ ................(1)\] When you solve equation (1) you will have the probability that A wins. Then you can find the required probability from: \[\large P(B \cup C)=1-P(A)\]

OpenStudy (anonymous):

thanks to you

OpenStudy (kropot72):

You're welcome :)